OpenStudy (anonymous):

the olivares family filled the gasoline tank of their minivan at the beginning of a long trip after driving 120 miles there were 16 gallons remaining in the tank if their car used gasoline at the rate of 1 gallon every 20 miles which equation describes the number of gallons y remaining in the tank after x miles?

OpenStudy (anonymous):

y=-1/20x+22 , y=20x+120 , y=20x-200 and y=1/20x-22